# Stabilizing and Solving Inverse Problems using Data and Machine Learning

### v1

@@ 4.3 Inverse Problem
$\boxed{\inf_{z\in Z}\frac{1}{2}\|u-\phi_{u,N,h}\circ\phi_{\text{de}}(z)\|_{H^{1}(\omega)}^{2}}$ (4.12)
@@ 5 Examples
The implementation used for the examples is based on the code presented in [28] which is publicly available at https://github.com/nmwsharp/neural-physics-subspaces. The GPU computations were performed on the Alvis cluster provided by NAISS (See Acknowledgements).
@@ 5.1 Linear Operator with Linear Data
We start with the fully linear case which we will build upon in the later examples. To construct a linear synthetic data set $\mathcal{G}$, we may pick a set of functions $\{\varphi_{j}\}_{j\in J}\subset H^{1/2}(\partial\Omega)$ and consider
@@
where $l$ is the circumference of $\Omega$ and $x$ is the counter-clockwise distance along the boundary starting from the point where the boundary crosses the first coordinate axis. We sample unperturbed coefficients $\widehat{g}_{j}\sim\mathcal{U}(-1,1)$ and perturbations $\delta_{j}\sim\mathcal{N}(0,0.0225)$. For each of the four meshes, we consider two values of the number of coefficients used to describe the boundary conditions; $N=9$ and $N=21$. We generate 1000 functions of the type (5.3) for each of the eight cases. Then, for every case, we compute a POD basis $\{\phi_{\mathrm{POD},j}\}_{j\in J}$ for the boundary using PCA on the data set. Unsurprisingly, the number of significant singular values turns out to be the number $N$ used in each case.
We use the POD boundary basis to compute an interior basis by solving Laplace’s equation with the finite element method (FEM). We take the discrete space $V_{h}$ to simply be the space of piecewise linear finite elements on the triangle mesh considered. The FEM interior basis $\{\phi_{\mathrm{FEM},j}\}_{j\in J}$ is computed by: For each $j\in J$, find $\phi_{\mathrm{FEM},j}\in V_{h}$ such that $\phi_{\mathrm{FEM},j}|_{\partial\Omega}=\phi_{\mathrm{POD},j}$ and
$(\nabla\phi_{\mathrm{FEM},j},\nabla v)_{\Omega}=0\quad\forall v\in V_{h}$ (5.4)
@@
We may now use the fact that Laplace’s equation is linear to superpose the FEM interior basis functions in a linear combination. This enables us to solve a linear inverse minimization problem over the coefficients in the linear combination. We present a demonstration of this process for the case with $N=9$ and the 82x82 mesh in Figure 4.
Figure 4: Optimization process for a linear inverse problem with noisy data using the FEM interior basis functions in Figure 3. The unperturbed data is shown in the first frame. The second frame is the same as the first but with added noise.
@@ 5.2 Nonlinear Operator with Linear Data
We again consider the linear data sets from the previous section, but here together with a nonlinear differential operator. Because of the nonlinearity, we cannot use the FEM interior basis and the superposition principle as in the fully linear case. Instead, we use a neural network to approximate the solution operator, i.e., the inverse of the nonlinear differential operator. The solution is still in the form of a finite element function, so the output of the network gives an approximation of the finite element solution. The input to the network is POD-coefficients corresponding to the same POD boundary basis functions as in the linear case in the previous section. We use the following nonlinear energy functional as the foundation for the loss function during training of the network.
[…]
@@
To measure the well-trainedness of the network, we, as an initial guiding measure, use the zero energy $E(u_{g=0})$, i.e., the value of the computed energy using the output from the network when an all zero vector is given as input. This, of course, corresponds to homogeneous Dirichlet boundary conditions and gives that the solution $u=0$ and thus that $E(u_{g=0})=0$. We also perform more rigorous studies of well-trainedness by computing the actual finite element solution with FEniCS [22] and comparing it to the network approximation. This is done by computing their average norm difference over 1000 problems, where for each problem we randomly select a coefficient vector with values from $\mathcal{N}(0,0.09)$. The difference is computed in both the $H_{0}^{1}$-norm ($H^{1}$-seminorm) and the $L^{2}$-norm. We also compute both the absolute and the relative norm differences, where the relative norm difference is the absolute difference divided by the norm of the finite element solution.
@@
The hyperparameter values for the problem sizes have been obtained by trial and error. In Table 2, we present training info for the four mesh sizes for coefficient vector length $N=9$ and $N=21$. The training has been performed on a single A100 GPU. For the largest mesh case (244x244), we have not been able to train with all elements present in the energy functional loss function (It has resulted in a NaN loss function value). To make it work, we have employed the trick of randomly selecting a fixed number of elements for every input vector during training, and only considering the energy functional contribution from those elements. The number of elements used is denoted “Els” in Table 2.
@@
Mesh Els Training time GPU Util Inference time $E(u_{g=0})$ $H_{0}^{1}$-error 1k-avg (rel) $L^{2}$-error 1k-avg (rel)
@@
Mesh Els Training time GPU Util Inference time $E(u_{g=0})$ $H_{0}^{1}$-error 1k-avg (rel) $L^{2}$-error 1k-avg (rel)
@@
With these neural networks we may solve the inverse minimization problem over the coefficient space. In Figure 5, a demonstration of this process is presented for the case of 21 input coefficients and the 244x244 mesh, i.e., the neural network whose training info is presented in the last row of Table 2(b).
Figure 5: Optimization process for a nonlinear inverse problem with noisy data using the operator network in the last row of Table 2(b) (21 input coefficients, 59049 output DoFs). The unperturbed data is shown in the first frame. The second frame is the same as the first but with added noise.
@@ 5.3 Nonlinear Operator with Nonlinear Data
We consider the same nonlinear differential operator with the same neural networks as in the previous section but here we add complexity by introducing an underlying nonlinear dependence on the input coefficients to the network. To construct such a nonlinear dependence we may pick a smooth function $a:X\rightarrow\mathbb{R}^{|J|}$, where $X$ is a parameter domain in $\mathbb{R}^{n_{X}}$ and consider boundary data of the form
$\mathcal{G}_{a}=\Bigl{\{}g_{i}=\sum_{j\in J}(a_{j}(x_{i})+\delta_{j})\phi_{\mathrm{POD},j}\,|\,i\in I\Bigr{\}}$ (5.6)
@@
We analyze data generated for the linear case with PCA and data generated for the quadratic case with both PCA and autoencoders. The results are shown in Figure 7.
Figure 7: Left: PCA results for both the linear and quadratic case. Right: Autoencoder results for the quadratic case.
@@
For the numerical results we take $\gamma=2$ and sample perturbations $\delta_{j}\sim\mathcal{N}(0,0.0225)$. We consider four cases:
@@
The coefficients, PCA results and autoencoder results are shown in Figure 8. For the autoencoder results we have used MLPs with 5 layers with the middle layer being the latent layer. The latent layer width has been varied and the hidden layer widths have all been fixed at 64. The activation function ELU has been applied to all layers except the last.
@@
Figure 8: Gaussian data examples. Left: Coefficients, unperturbed (blue) and perturbed (red). Middle: PCA result. Right: Autoencoder result.
@@
From the Gaussian data examples presented in Figure 8 we note something interesting. If the number of bell curves $L$ is divisible by the latent dimension $n_{x}$, the PCA suggests that the underlying structure has dimension $L$. If $L$ is not divisible by $n_{x}$, the PCA instead suggests that this dimension is $n_{x}L$. In the third example with results presented in Figure 8(c), we have $(n_{x},L)=(3,7)$. Here the PCA suggests that the underlying dimension is 21, whereas the corresponding autoencoder study suggests that a reduction down to 9 dimensions could provide the same improvement as a reduction down to 17.
@@
In light of the above, we may take the autoencoder with latent layer width = 9 from this case and connect its decoder to the input of the operator network for the 244x244 mesh with 21 input coefficients. We may thus solve the inverse minimization problem over a 9 dimensional latent space instead of a 21 dimensional coefficient space. We present a demonstration of this process in Figure 9.
@@
Figure 9: Optimization process for a nonlinear inverse problem with noisy data, again using the operator network in the last row of Table 2(b) (21 input coefficients, 59049 output DoFs) but here together with the decoder from the right frame in Figure 8(c) that maps from a 9-dimensional latent space to a 21 dimensional coefficient space. The unperturbed data is shown in the first frame. The second frame is the same as the first but with added noise.

### v2

@@ 4.3 Inverse Problem
$\boxed{\inf_{z\in Z}\frac{1}{2}\|u_{0}-(\phi_{u,N,h}\circ\phi_{\text{de}})(z)\|_{L^{2}(\omega)}^{2}}$ (4.12)
@@ 5 Examples
The geometry is the same in all the examples. We take the solution domain $\Omega:=(-0.5,0.5)^{2}\subset\mathbb{R}^{2}$ and the reference domain $\omega\subset\Omega$ to be the u-shaped domain defined by
$\omega:=\{(x,y)\in\mathbb{R}^{2}\,|\,x<0.25\land(x<-0.25\lor y<-0.25\lor y>0.25)\}$ (5.1)
see Figure 3.
Figure 3: Domain used in all numerical examples with the subdomain $\omega$ indicated.
When solving the inverse problems in practice, we use data $u_{0}\in V_{h}$. We also minimize the mean squared error (MSE) over the DoFs belonging to $\omega$ instead of the squared $L^{2}(\omega)$ norm of the error, which is valid since they are equivalent on $V_{h}$ from the Rayleigh quotient. The only “stabilization” we use for the inverse problem is that the boundary data is finite-dimensional and that this dimension together with the mesh size $h$ both are small enough. We point out that no additional stabilization, such as including penalty terms, is used. The criterion for when a minimization process is considered to have converged is based on the change of significant digits of the MSE. For the fully linear problem we consider it converged when at least three significant digits remain constant, and for all the nonlinear problems when at least two significant digits remain constant. This is in turn based on when both the optimization variables and the visual representation of the output do not seem to change anymore and has been obtained by testing.
The implementation used for the examples is based on the code presented in [38] which is publicly available at https://github.com/nmwsharp/neural-physics-subspaces. All inverse problem minimizations have been performed with the Adam optimizer with a step size = 0.1 on an Apple M1 CPU. The GPU computations were performed on the Alvis cluster provided by NAISS (See Acknowledgements).
@@ 5.1 Linear Operator with Linear Data
We start with the fully linear case which we will build upon in the later examples. To construct a linear synthetic data set $\mathcal{G}$, we may pick a set of functions $\{\varphi_{j}\}_{j\in J}\subset H^{1/2}(\partial\Omega)$ where $J$ is some index set, and consider
@@
where $l$ is the circumference of $\Omega$ and $x$ is the counter-clockwise distance along the boundary starting from the point where the boundary crosses the first coordinate axis. We sample unperturbed coefficients $\widehat{g}_{j}\sim\mathcal{U}(-1,1)$ and perturbations $\delta_{j}\sim\mathcal{N}(0,0.0225)$. For each of the four meshes, we consider two values of the number of coefficients used to describe the boundary conditions; $N=9$ and $N=21$. We generate 1000 functions of the type (5.4) for each of the eight cases. Then, for every case, we compute a POD basis $\{\varphi_{\mathrm{POD},j}\}_{j\in J}$ for the boundary using PCA on the data set. Unsurprisingly, the number of significant singular values turns out to be the number $N$ used in each case.
We use the truncated POD boundary basis corresponding to significant singular values to compute and interior basis. We do this by solving Laplace’s equation with the finite element method (FEM). We take the discrete space $V_{h}$ to simply be the space of piecewise linear finite elements on the triangle mesh considered. The FEM interior basis $\{\varphi_{\mathrm{FEM},n}\}_{n=1}^{N}$ is computed by: For each $n=1,...,N$, find $\varphi_{\mathrm{FEM},n}\in V_{h}$ such that $\varphi_{\mathrm{FEM},n}|_{\partial\Omega}=\varphi_{\mathrm{POD},n}$ and
$(\nabla\varphi_{\mathrm{FEM},n},\nabla v)_{\Omega}=0\quad\forall v\in V_{h}$ (5.5)
@@
We may now use the fact that Laplace’s equation is linear to superpose the FEM interior basis functions in a linear combination.
$u_{h,\text{lin},N}=\sum_{n=1}^{N}c_{n}\varphi_{\mathrm{FEM},n}$ (5.6)
This enables us to solve a linear inverse minimization problem over the coefficients $(c_{1},...,c_{N})\in\mathbb{R}^{N}$ in the linear combination. We present a demonstration of this process for the case with $N=9$ and the 82x82 mesh in Figure 5. There it can clearly be observed that the finite element solution given by the linear combination approaches the noisy data and the reference solution as the optimization progresses.
Figure 5: Optimization process over a 9-dimensional coefficient space for a linear inverse problem with noisy data. Here, the FEM interior basis functions in Figure 4 are used. The unperturbed data is shown in the first frame. The second frame is the same as the first but with added noise sampled from $\mathcal{U}(-0.05,0.05)$. The last frame shows $\omega$ and the reference solution used for the data which was obtained by taking $c_{1}=10,c_{9}=3.023,$ and all other $c_{n}$’s = 0 in (5.6). The penultimate frame shows the optimization’s MSE-converged reconstruction of the reference solution. The MSE converged after 861 iterations with the Adam optimizer with a step size = 0.1. This took 10.3 s on an Apple M1 CPU. The MSE’s between the reference solution and the converged reconstruction are: on $\omega$ (used in optimization), $\text{MSE}_{\omega}$ = 8.28e-4; on the convex hull of $\omega$, $\text{MSE}_{\text{co}(\omega)}$ = 9.12e-4; and on its complement, $\text{MSE}_{\text{co}(\omega)^{c}}$ = 2.66e-3.
@@ 5.2 Nonlinear Operator with Linear Data
We again consider the linear data sets from the previous section, but here together with a nonlinear differential operator. Because of the nonlinearity, we cannot use the FEM interior basis and the superposition principle as in the fully linear case. Instead, we use a neural network to approximate the solution operator, i.e., the inverse of the nonlinear differential operator. The solution is still in the form of a finite element function, so the output of the network gives an approximation of the finite element solution. The input to the network is the POD coefficients $(p_{1},...,p_{N})$ corresponding to the same significant POD boundary basis functions as in the fully linear case. We use the following nonlinear energy functional as the foundation for the loss function during training of the network.
$E(v)=\int_{\Omega}\frac{1}{2}(1+v^{2})|\nabla v|^{2}\mathop{}\!\mathrm{d}x$ (5.7)
@@
To measure the well-trainedness of the network, we, as an initial guiding measure, use the zero energy $E(\phi_{u,N,h}(0))$, i.e., the value of the computed energy using the output from the network when an all zero vector is given as input. This, of course, corresponds to homogeneous Dirichlet boundary conditions and gives that the solution $u=0$ and thus that $E(0)=0$. We also perform more rigorous studies of well-trainedness by computing the actual finite element solution with FEniCS [29] and comparing it to the network approximation. This is done by computing their average norm difference over 1000 problems, where for each problem we randomly select a coefficient vector with values from $\mathcal{N}(0,0.09)$. The difference is computed in both the $H_{0}^{1}$-norm ($H^{1}$-seminorm) and the $L^{2}$-norm. We also compute both the absolute and the relative norm differences, where the relative norm difference is the absolute difference divided by the norm of the finite element solution.
@@
The hyperparameter values for the problem sizes have been obtained by trial and error. In Table 2, we present training info for the four mesh sizes for coefficient vector length $N=9$ and $N=21$. The training has been performed on a single A100 GPU. For the largest mesh case (244x244), we have not been able to train with all elements present in the energy functional loss function (It has resulted in a NaN loss function value). To make it work, we have employed the trick of randomly selecting a fixed number of elements for every input vector during training, and only considering the energy functional contribution from those elements. The number of elements used is denoted “Els” in Table 2. It can be observed from the zero energies and norm errors in Table 2 that the operator networks generally become more accurate with finer meshes if all elements are used in the energy computation. In both cases with the 244x244 mesh, the trend in higher accuracy is broken. This is reasonable since only a few elements, instead of all, are used in the energy computation for the loss function.
@@
Mesh Els Training time GPU Util Inference time $E(\phi_{u,N,h}(0))$ $H_{0}^{1}$-error 1k-avg (rel) $L^{2}$-error 1k-avg (rel)
@@
Mesh Els Training time GPU Util Inference time $E(\phi_{u,N,h}(0))$ $H_{0}^{1}$-error 1k-avg (rel) $L^{2}$-error 1k-avg (rel)
@@
With these neural networks we may solve the inverse minimization problem over the coefficient space. In Figure 6, a demonstration of this process is presented for the case of 21 input coefficients and the 244x244 mesh, i.e., the neural network whose training info is presented in the last row of Table 2(b). In Figure 6 it can clearly be observed that the approximate finite element solution given by the operator network approaches the noisy data and the reference solution as the optimization progresses.
Figure 6: Optimization process over a 21-dimensional coefficient space for a nonlinear inverse problem with noisy data. Here, the operator network in the last row of Table 2(b) (21 input coefficients, 59049 output DoFs) is used. The unperturbed data is shown in the first frame. The second frame is the same as the first but with added noise sampled from $\mathcal{U}(-0.05,0.05)$. The last frame shows $\omega$ and the reference solution used for the data which was obtained by taking $p_{14}=10$ and all other $p_{n}$’s = 0. The penultimate frame shows the optimization’s MSE-converged reconstruction of the reference solution. The MSE converged after 2843 iterations with the Adam optimizer with a step size = 0.1. This took 140.2 s on an Apple M1 CPU. The MSE’s between the reference solution and the converged reconstruction are: on $\omega$ (used in optimization), $\text{MSE}_{\omega}$ = 8.36e-4; on the convex hull of $\omega$, $\text{MSE}_{\text{co}(\omega)}$ = 8.38e-4; and on its complement, $\text{MSE}_{\text{co}(\omega)^{c}}$ = 1.91e-3.
@@ 5.3 Nonlinear Operator with Nonlinear Data
We consider the same nonlinear differential operator with the same neural networks as in the previous section but here we add complexity by introducing an underlying nonlinear dependence on the input coefficients to the network. To construct such a nonlinear dependence we may pick a smooth function $a:X\rightarrow\mathbb{R}^{|J|}$, where $X$ is a parameter domain in $\mathbb{R}^{n_{X}}$, and for some index set $I$, consider boundary data of the form
$\mathcal{G}_{a}=\Bigl{\{}g_{i}=\sum_{j\in J}(a_{j}(x_{i})+\delta_{j})\varphi_{\mathrm{POD},j}\,|\,i\in I\Bigr{\}}$ (5.8)
@@
We analyze data generated for the linear case with PCA and data generated for the quadratic case with both PCA and autoencoders. For the autoencoders we have used MLPs with 6 layers (5 hidden, 1 output) with the third layer being the latent layer. The latent layer width has been varied and the remaining hidden layer widths have all been fixed at 64. The activation function ELU has been applied to all layers except the last. The training has been performed with the Adam optimizer exactly as for the operator networks, i.e., $10^{6}$ iterations with a decreasing learning rate. The batch size has been 64. The training time for a single autoencoder (fixed latent layer width) on an Apple M1 CPU has typically been in the range 240 – 270 s.
The results from both the PCA and autoencoder analysis are presented in Figure 8. The PCA results give a 3-dimensional latent space in the linear case and a 6-dimensional in the quadratic. This is evident from the number of significant singular values for the different cases. The autoencoder results suggest the existence of both a 3- and a 6-dimensional latent space in the quadratic case. This can be deduced from the two plateaus for the two perturbed cases: one at latent layer widths 3 – 5 and one at 6 – 8. The autoencoders thus manage to find the underlying 3-dimensional structure in the quadratic case whereas PCA does not.
Figure 8: Left: PCA results for both the linear and quadratic case. The plots show the singular values of the coefficients in decreasing order for three different perturbations: unperturbed and two perturbed (standard deviation = 0.5 and 1). Right: Autoencoder results for the quadratic case for the same three perturbations as PCA but used during training. The plots show the average mean squared reconstruction error over unperturbed test data for different latent layer widths both on a linear and logarithmic scale.
@@
For the numerical results we take $\gamma=2$ and sample perturbations $\delta_{j}\sim\mathcal{N}(0,0.0225)$ (standard deviation = 0.15). We consider four cases:
@@
We analyze data generated for these cases with both PCA and autoencoders. For the autoencoders we have used MLPs with 5 layers (4 hidden, 1 output) with the middle layer being the latent layer. The latent layer width has been varied and the remaining hidden layer widths have all been fixed at 64. The activation function ELU has been applied to all layers except the last. The training has been performed with the Adam optimizer exactly as for the operator networks, i.e., $10^{6}$ iterations with a decreasing learning rate. The batch size has been 64. The training time for a single autoencoder (fixed latent layer width) on an Apple M1 CPU has typically been in the range 210 – 250 s.
The bell curves for the coefficients, PCA results and autoencoder results are presented in Figure 9. The PCA results show something interesting. If the number of bell curves $L$ is divisible by the latent dimension $n_{X}$, PCA gives that the underlying structure has dimension $L$. If $L$ is not divisible by $n_{X}$, PCA instead gives that this dimension is $n_{X}L$. For example, for $(n_{X},L)=(2,5)$ in Figure 9(a), PCA gives latent dimension = 10, and for $(n_{X},L)=(3,6)$ in Figure 9(b), PCA gives latent dimension = 6. This phenomenon is easily understood by the number of unique combinations of latent parameters $x_{k}$ and bell curves, characterized by $x_{0,l}$, in the construction of the coefficients given by (5.10). The autoencoder results all suggest the existence of latent spaces of a lower dimension than given by PCA. This is most clearly seen from the existence of plateaus for the two perturbed cases (standard deviation = 0.075 and 0.15) on the logarithmic scale in all four cases. However, the suggested latent dimension does match the actual one as well as in the previous example with polynomial data, hinting at the higher complexity of the Gaussian data. This is especially true in the cases where $n_{X}$ does not divide $L$.
@@
Figure 9: Gaussian data examples. Left: Bell curves used for the coefficients, unperturbed (blue) and perturbed (red). Middle: PCA results for unperturbed data (blue) and perturbed (red, standard deviation = 0.15). The plots show the singular values of the coefficients in decreasing order. Right: Autoencoder results for three different perturbations used during training: unperturbed and two perturbed (standard deviation = 0.075 and 0.15). The plots show the average mean squared reconstruction error over unperturbed test data for different latent layer widths both on a linear and logarithmic scale.
@@
In the third Gaussian data example with results presented in Figure 9(c), we have $(n_{X},L)=(3,7)$. Here the PCA suggests that the underlying dimension is 21 (number of significant singular values), whereas the corresponding autoencoder study suggests that a reduction down to 9 dimensions could provide the same improvement as a reduction down to 17 in the case of the autoencoders trained on perturbed data (9 and 17 give roughly the same error). In light of the above, we may take an autoencoder with latent layer width = 9 from this case and connect its decoder to the input of the operator network for the 244x244 mesh with 21 input coefficients. We may thus solve an inverse minimization problem over a 9-dimensional latent space instead of a 21-dimensional coefficient space. We present demonstrations of this process in Figures 1012. A summary of the optimization results for these three demonstrations and also the one in Figure 6 is given in Table 3.
The main difference between the three demonstrations is the decoder used. First in Figures 1011, we use the decoder from the “sd = 0” autoencoder, meaning it was trained on unperturbed data. The first of these two demonstrations is for clean data, $u_{0}$ in $\omega$, and the second for noisy. We see that the two optimization processes are essentially the same but find it instructive to present both as the clean data case functions as a reference. Second, in Figure 12, we use the decoder from the “sd = 0.15” autoencoder, meaning it was trained on perturbed data with perturbations from $\mathcal{N}(0,0.0225)$. From the logarithmic scale in the right frame in Figure 9(c) we see that the reconstruction errors of the two autoencoders differ substantially, by several orders of magnitude. Comparing the corresponding optimization processes, we also see that using the “sd = 0” decoder (Figure 11) produces a much more accurate reconstruction compared to the “sd = 0.15” decoder (Figure 12) that fails to do so.
The reconstructions in all three decoder cases, and especially the last, are less accurate compared to the case with only the operator network presented in Figure 6, as can be seen from both the figures and the MSE’s in Table 3. This is reasonable since the reference solution in all four cases is the same network output corresponding to a specific coefficient input and in the case with only the operator network we optimize in this coefficient space whereas in the decoder cases in some latent space. It is simply not guaranteed that the decoders may attain this specific coefficient input when mapping from the latent space. One reason being that a single change in any of the 9 latent variables can affect all the 21 coefficients. Comparing the MSE’s on the different subdomains in Table 3, we see that in all four cases it is smaller on the convex hull of $\omega$ than on the complement as expected. This is also true for the fully linear case (corresponding results are presented in the caption of Figure 5). The average iteration times presented in Table 3 are essentially the same for the four cases. Something that is positive for using decoders, but maybe not so surprising considering how much smaller the decoder MLP’s are in comparison to the operator MLP. In summary, autoencoders may be used to reduce the dimension of the optimization space (latent instead of coefficient space), but to really gain from such a reduction and to maintain accuracy, care needs to be taken in how the reduction mapping is constructed. We point out that the MLP approach considered here is rather simple and that we believe there is substantial room for improvement by considering more sophisticated methods.
As final remarks we point out that taking some output of the method under consideration as the reference solution, as is done here, is typically not a proper choice since it is too idealized. However, here we make this choice to put more focus on the effects of latent space optimization. We also point out that all the optimization processes involving neural networks presented here have been for the rougher networks: the operator network in the last row of Table 2(b) and the autoencoders in Figure 9(c) have alternatives with better measures of well-trainedness. The idea behind this being that if the concept works to some degree in the harder cases, it should work even better in the easier ones.
@@
Figure 10: Optimization process over a 9-dimensional latent space for a nonlinear inverse problem with clean data. Again, the operator network in the last row of Table 2(b) (21 input coefficients, 59049 output DoFs) is used, but here together with the “sd = 0” decoder from the right frame in Figure 9(c). The decoder maps from a 9-dimensional latent space to a 21-dimensional coefficient space. The last frame shows $\omega$ and the reference solution used for the data which was obtained by taking $p_{14}=10$ and all other $p_{n}$’s = 0. The penultimate frame shows the optimization’s MSE-converged reconstruction of the reference solution. The MSE converged after 1481 iterations with the Adam optimizer with a step size = 0.1. This took 68.9 s on an Apple M1 CPU. The MSE’s between the reference solution and the converged reconstruction are: on $\omega$ (used in optimization), $\text{MSE}_{\omega}$ = 2.22e-3; on the convex hull of $\omega$, $\text{MSE}_{\text{co}(\omega)}$ = 1.52e-3; and on its complement, $\text{MSE}_{\text{co}(\omega)^{c}}$ = 4.83e-3.
Figure 11: Optimization process over a 9-dimensional latent space for a nonlinear inverse problem with noisy data. Again, the operator network in the last row of Table 2(b) (21 input coefficients, 59049 output DoFs) is used together with the “sd = 0” decoder from the right frame in Figure 9(c). The decoder maps from a 9-dimensional latent space to a 21-dimensional coefficient space. The unperturbed data is shown in the first frame. The second frame is the same as the first but with added noise sampled from $\mathcal{U}(-0.05,0.05)$. The last frame shows $\omega$ and the reference solution used for the data which was obtained by taking $p_{14}=10$ and all other $p_{n}$’s = 0. The penultimate frame shows the optimization’s MSE-converged reconstruction of the reference solution. The MSE converged after 1018 iterations with the Adam optimizer with a step size = 0.1. This took 47.5 s on an Apple M1 CPU. The MSE’s between the reference solution and the converged reconstruction are: on $\omega$ (used in optimization), $\text{MSE}_{\omega}$ = 3.07e-3; on the convex hull of $\omega$, $\text{MSE}_{\text{co}(\omega)}$ = 2.37e-3; and on its complement, $\text{MSE}_{\text{co}(\omega)^{c}}$ = 5.25e-3.
Figure 12: Optimization process over a 9-dimensional latent space for a nonlinear inverse problem with noisy data. Again, the operator network in the last row of Table 2(b) (21 input coefficients, 59049 output DoFs) is used, but here together with the “sd = 0.15” decoder from the right frame in Figure 9(c). The decoder maps from a 9-dimensional latent space to a 21-dimensional coefficient space. The unperturbed data is shown in the first frame. The second frame is the same as the first but with added noise sampled from $\mathcal{U}(-0.05,0.05)$. The last frame shows $\omega$ and the reference solution used for the data which was obtained by taking $p_{14}=10$ and all other $p_{n}$’s = 0. The penultimate frame shows the optimization’s MSE-converged reconstruction of the reference solution. The MSE converged after 212 iterations with the Adam optimizer with a step size = 0.1. This took 11.0 s on an Apple M1 CPU. The MSE’s between the reference solution and the converged reconstruction are: on $\omega$ (used in optimization), $\text{MSE}_{\omega}$ = 1.99e-2; on the convex hull of $\omega$, $\text{MSE}_{\text{co}(\omega)}$ = 1.44e-2; and on its complement, $\text{MSE}_{\text{co}(\omega)^{c}}$ = 5.51e-2.
Table 3: Summary of optimization results for all nonlinear inverse problems using an operator network. All problems have the same reference solution and use the operator network in the last row of Table 2(b) (21 input coefficients, 59049 output DoFs). The optimization processes for the problems are presented in Figures 6, 1012, respectively. In the table, “Op” means the operator network, “dec” means decoder, “sd = x” means what perturbation was added to the training data for the decoder, and “$\delta_{\omega}$” means that noisy data was used for the inverse problem.
Configuration Iterations Avg iter time $\text{MSE}_{\omega}$ $\text{MSE}_{\text{co}(\omega)}$ $\text{MSE}_{\text{co}(\omega)^{c}}$
Op, $\delta_{\omega}$ 2843 4.93e-2 s 8.36e-4 8.38e-4 1.91e-3
Op + dec “sd = 0” 1481 4.65e-2 s 2.22e-3 1.52e-3 4.83e-3
Op + dec “sd = 0”, $\delta_{\omega}$ 1018 4.67e-2 s 3.07e-3 2.37e-3 5.25e-3
Op + dec “sd = 0.15”, $\delta_{\omega}$ 212 5.19e-2 s 1.99e-2 1.44e-2 5.51e-2